4. Given that HCF (306, 657) = 9, find LCM (306, 657).

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
HCF×LCM=Product of the two given numbers Therefore, 9 × LCM = 306 × 657 LCM = (306×657)/9 = 22338 Hence, LCM(306,657) = 22338